OpenStudy (anonymous):

solve for x show all work -2x+5=4x-7

OpenStudy (samwisegamge):

do you have a ti84 calculator?

OpenStudy (kittiwitti1):

Isolate the x Here, I'll start out: \[-2x+5=4x-7\]\[-2x+5-5=4x-7-5\]\[-2x=4x-2\]\[-2x-4x=4x-2-4x...\]

OpenStudy (kittiwitti1):

Can you do it from here or should I keep going? :-P

OpenStudy (anonymous):

can you keep going please im a little confused

OpenStudy (kittiwitti1):

http://prntscr.com/aa33pe

OpenStudy (kittiwitti1):

still stuck?

OpenStudy (anonymous):

no thank you so much

OpenStudy (kittiwitti1):

alright np :-)

OpenStudy (kittiwitti1):

medals are given by clicking "best response", the blue button :-)

OpenStudy (kittiwitti1):

^-^